find the Taylor Series explanaion for f(x) = x^-2 around the point a=2. what is the radius od convergence?
Answer #1

solution is given that finde 2-2 frasat ,a=2 22 f(2)= - 22 23 {m>- they +|) - put n=2 +(2) = +121= t(2) SEP t (n) z -29Taylor Schips at naa - ہ رارار-2) - ردلار (-) + رو) - (۶) a=2 put S یہ رہ کر( : ( f ( x رع-7) + رد/ ر2 -9 ) + ر )+ 2 با ر یدر9373 let an= (-18n /n+)/2-223 2372 Teske ant) =(-1.5/3+2)/2+2 3+) By Ratio Test lim 1 ant), A 102337) (+3)(3-2)}) 2342 (-)/